40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6620     Entry time: Tue May 8 03:33:24 2012
Author: Jenne 
Type: Update 
Category: SUS 
Subject: New misalign / restore scripts for IFO_ALIGN screen 

Since Den wasn't able to fix c1sus (make it talk to the framebuilder) before he left a few hours ago, I decided to do some housekeeping rather than actual locking. 

I wrote new save / misalign / restore scripts for all of the suspended optics on the C1IFO_ALIGN screen.  Adding save / restore versions for the mode cleaner optics should be quick and easy.  Now when you use the ! button for each optic, it points you to the new scripts.  I still have the burt capabilities there, but the restore script has the burt-restore line commented out.


SAVE:  burt-save the PIT_COMM and YAW_COMM values, as well as write those values and the date to a text file.

MISALIGN:  Turn off oplevs, move 100 steps of 0.01 in the "+" direction.

RESTORE: Move ~100 steps toward the saved value, until you're within 0.001 of the saved value (step size is "saved val" minus "current val" divided by 100).  Then just write the saved value to the slider (otherwise if the slider were touched between the last "save" and the restore, we might not be able to step precisely to the value we want). Turn oplevs back on.

Scripts are in the same place the old ones used to live:  ...../caltech/c1/medm/c1ifo/cmd/   New scripts are C1IFO_OPTIC(save/restore/misalign)_soft.cmd

I'm checking this one off of the to-do list.


Good things:  (a) I remembered / re-learned / just plain learned a lot about scripting.  (b) the optics are now walked slowly over to their misaligned state, and slowly walked back.  The past regime had the optics suddenly kicked over by a lot, sometimes enough to trip / come close to tripping watchdogs, which was never good.

Bad things: it took a long time.  Now it's bedtime.

ELOG V3.1.3-